OpenStudy (anonymous):

(1+sinx)(1-sinx)=cos^2 x

OpenStudy (anonymous):

\[(1-\sin(x))(1+\sin(x))=1-\sin^2(x)=1-(1-\cos^2(x))=\cos^2(x)\]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

welcome